What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0?
A.) 4x - y = 13
B.) 4x + y = 13
C.) 4x + y = -13

Answer by ikleyn(52797)

.
What is the equation of the line, in standard form, that passes through (4, -3) and is parallel to the line whose equation is 4x + y - 2 = 0?
A.) 4x - y = 13
B.) 4x + y = 13
C.) 4x + y = -13
~~~~~~~~~~~~~~~~~~

First of all, the line parallel to the given line 4x + y - 2 = 0 has the form 

4x + y -c = 0,      (1)

where "c" is some constant value.

To find this constant value "c", use the condition that the line passes through the point (4,-3) and substitute x= 4, y= -3 into equation (1).
You will get

4*4 + (-3) -c = 0,   or   16-3 = c,   or  c = 13.

Thus the equation of the line under the question is 

4x + y - 13 = 0,  or, which is the same,   4x + y = 13.


Answer. Option C).

Solved.